# Enabling Robust Exoplanet Atmospheric Retrievals with Gaussian Processes

### v1

@@ Abstract
Atmospheric retrievals are essential tools for interpreting exoplanet transmission and eclipse spectra, enabling quantitative constraints on the chemical composition, aerosol properties, and thermal structure of planetary atmospheres. The James Webb Space Telescope (JWST) offers unprecedented spectral precision, resolution, and wavelength coverage, unlocking transformative insights into the formation, evolution, climate, and potential habitability of planetary systems. However, this opportunity is accompanied by challenges: modeling assumptions and unaccounted-for noise or signal sources can bias retrieval outcomes and their interpretation. To address these limitations, we introduce a Gaussian Process (GP)-aided atmospheric retrieval framework that flexibly accounts for unmodeled features in exoplanet spectra, whether global or localized. We validate this method on synthetic JWST observations and show that GP-aided retrievals reduce bias in inferred abundances and better capture model-data mismatches than traditional approaches. We also introduce the concept of mean squared error to quantify the trade-off between bias and variance, arguing that this metric more accurately reflects retrieval performance than bias alone. We then reanalyze the NIRISS/SOSS JWST transmission spectrum of WASP-96 b, finding that GP-aided retrievals yield broader constraints on CO2 and H2O, alleviating tension between previous retrieval results and equilibrium predictions. Our GP framework provides precise and accurate constraints while highlighting regions where models fail to explain the data. As JWST matures and future facilities come online, a deeper understanding of the limitations of both data and models will be essential, and GP-enabled retrievals like the one presented here offer a principled path forward.
@@ 2.1 Forward Model
The vertical temperature structure is parameterized following the prescription of Madhusudhan & Seager (2009a). Inhomogeneous clouds and hazes (aerosols) are modeled using an adaptation of the single-sector model from Welbanks & Madhusudhan (2021) where instead of a cloud-top pressure, we consider a grey cloud with an opacity of . The contribution from hazes is modeled as a modification to Rayleigh-scattering with a slope of and enhancement . The inhomogeneous cloud/haze cover is modeled as a linear combination of a clear and cloudy/hazy atmosphere following Line & Parmentier (2016) where is a free parameter describing the cloud/haze fraction. Finally, we include a scaling parameter, R, that accounts for uncertainty in the reported radius of the planet and its degeneracy with the reference pressure (Welbanks & Madhusudhan, 2019). The spectral model is evaluated at a resolution of before being binned to the resolution of the observations. The Bayesian inference is performed using nested sampling algorithm (Feroz et al., 2009) via PyMultiNest (Buchner et al., 2014).
@@ 2.2 Retrieval Framework
where n is the number of wavelength bins, is the observed flux (or transit depth) in the bin with associated uncertainty , and is the model prediction at that wavelength.
@@
To account for these correlations, we include a Gaussian process (GP) in our retrieval framework (Figure 1). A GP is defined by two components: a mean function, , which in our case corresponds to the modeled atmospheric spectrum, and a covariance matrix, , which captures correlations between data points and is constructed from one or more parametric kernel functions, , each with associated hyperparameters (Rasmussen & Williams, 2005). Each element quantifies the covariance between data points and , while the diagonal elements represent the white noise variance at each wavelength bin. When noise is uncorrelated and well-characterized, the off-diagonal terms vanish, and the standard Gaussian likelihood is recovered. A global kernel captures long-range correlations across the entire spectrum, whereas local kernels can be used to model confined regions exhibiting strongly correlated residuals. Figure 2 shows a visual example of a covariance matrix that includes both global and local kernels.
@@
where is the residual between the observed spectrum and the binned model evaluation, and is the GP covariance matrix. In the traditional retrieval case, where data points are assumed to be uncorrelated, reduces to a diagonal matrix with elements (see Figure 2). Under this assumption, Equation 3 simplifies to the standard one-dimensional form shown in Equation 2. Because the GP acts directly on the residuals, it is agnostic to whether discrepancies arise from imperfections in the data or from inadequacies in the model. This allows the retrieval to account for correlated structure in the residuals without requiring explicit identification of its source.
Here, we adopt a covariance matrix composed of both global and local kernels. For the global kernel, we assume a single amplitude and correlation length scale applied uniformly across the spectrum. Local kernels, in contrast, are fitted to capture correlated structure confined to specific regions. These localized deviations can arise from unmodeled spectral features such as missing absorbers or contamination in the data (e.g., Czekala et al., 2015), as well as from artifacts introduced during data reduction (e.g., Radica et al., 2023; Holmberg & Madhusudhan, 2023). Accounting for these features with local kernels allows the retrieval to marginalize over regional model–data mismatches that might otherwise bias inferred atmospheric properties. Notably, such biases may not be evident in traditional retrieval frameworks, as residuals in one part of the spectrum can still bias the retrieval outcome for absorbers with features elsewhere in the spectrum (see e.g., Welbanks et al., 2023). In addition, the retrieved hyperparameters of the local kernels can help identify which regions of the spectrum contribute most to poor fits and quantify the extent of the correlated deviations.
@@
where represents the traditional diagonal covariance matrix, is the global kernel, and is the local kernel.
@@ 2.3 Kernel Selection
The global kernel is defined by a characteristic amplitude and length scale , where the length scale represents the distance over which wavelength bins exhibit non-negligible correlation (Rasmussen & Williams, 2005). Typically, a positive semi-definite kernel, i.e., one for which for all , is applied across the entire mean function. In this work, we model the global covariance using a Gaussian (radial basis function; RBF) kernel,
where is the distance between any two wavelength bins and . Other studies, including Czekala et al. (2015) and Meech et al. (2022), have used a Matérn-3/2 kernel instead, which can be considered a rougher variant of the RBF kernel. We compare the two on the synthetic spectrum in Section 3 and find little difference in performance; however, the RBF kernel yields a marginally more accurate H2O abundance, and so we adopt it throughout this work. In addition, we focus specifically on observations obtained with NIRISS/SOSS, which has a point spread function (PSF) oversampled by 2–3 pixels along the dispersion axis (Albert et al., 2023). Under these conditions, a Gaussian approximation is well-suited to model the covariance between adjacent pixels.
In addition to a global covariance, localized regions of high residuals can be modeled using local kernels. We assume the local kernel to be a Gaussian function (RBF) with amplitude , length scale , and central wavelength , such that
The RBF kernel acts as a smooth, low-resolution approximation to unresolved or unmodeled spectral features. While it can effectively account for unknown absorbers or contamination, its functional form is not flexible enough to mimic high-resolution line shapes. This makes it unlikely to compete with physically motivated opacity sources when included in the model, ensuring that the GP captures residual structure without affecting the detection or characterization of real atmospheric features.
@@ 3 Validation on Synthetic NIRISS/SOSS Spectra
The resulting atmospheric model was then binned to the wavelengths and widths from Radica et al. (2023). In addition to white noise drawn from the reported precision of the JWST/NIRISS SOSS spectrum in Radica et al. (2023), we inject global correlated noise across the entire wavelength range, with an amplitude of ppm and a correlation length scale of m. We also inject a localized high-correlation feature centered at 1.65 m. The final synthetic spectrum is generated as a single random draw from a multivariate normal distribution, incorporating both white and correlated noise. This approach simulates realistic observational scatter while preserving the imposed correlation structure. In doing so, we test whether the retrieved hyperparameters can inform inferences about the underlying correlation structure and the specific wavelengths most affected. The model, injected noise, and resulting synthetic data are shown in Figure 3. In addition to the noticeable feature at 1.65 m, the simulated noise results in a generally higher transit depth relative to the input model at wavelengths m.
@@
which corresponds to the specific case of Equation 4 with one local kernel. Table 1 provides the priors for the parameters and hyperparameters considered.
@@ 3.1 Retrieval Performance on Synthetic WASP-96 b Data
Although the traditional retrieval produces the most precise posterior distributions, it yields highly inaccurate results due to the injected correlated noise. The median retrieved abundances of H2O, CO2, and CH4 are offset from the true values by 4.2, 6.3, and 13.6, respectively (Figure 4). The latter corresponds to a vanishingly small probability of agreement under standard Gaussian assumptions, on the order of , illustrating the extreme overconfidence of a retrieval framework that neglects correlated noise.
@@
In the “local-only” retrieval, a single local kernel is used to capture a high-correlation feature in the spectrum. This configuration enables the retrieval to constrain the injected feature’s amplitude and central wavelength, but it does not account for the global correlations present across the spectrum. As a result, the retrieval remains biased in key atmospheric parameters: the H2O and CO2 abundances are still offset by 1.5 and 1.6, respectively. However, the local kernel successfully mitigates the effect of the injected feature near 1.7 m, which overlaps with a CH4 absorption band, resulting in a CH4 abundance much closer to the true value (0.4 offset). This suggests that while local kernels can improve model–data agreement in targeted regions, they are insufficient for capturing broader correlated structure.
@@
In contrast to the traditional case, the retrieval using a global kernel recovers all gas abundances to within 2 of their true values, but at the expense of precision. For example, the CO2 abundance is now much more accurate (0.2 offset compared to 6.3 in the traditional case), but the 1 posterior interval spans nearly 4 dex (Figure 4). As a result, the retrieval provides only weak constraints that are insufficient to inform parameters such as atmospheric metallicity or the carbon-to-oxygen ratio.
@@
Finally, we consider the retrieval that includes both global and local kernels. This configuration enables the GP to constrain all hyperparameters to within 2 of their injected values (Figure 4, bottom), including successful identification of the injected feature at 1.65 m. It also yields atmospheric parameters that are much closer to the true values, without substantially increasing the posterior variance. The retrieved median abundances of H2O, CO2, and CH4 differ from their true values by only 0.28, 0.26, and 0.10, respectively (Figure 4). The Na abundance is measured with higher precision than in the global-only case and greater accuracy than in the traditional and local-only retrievals, although it is still overestimated relative to the input. This overestimation may be partially attributable to the specific noise realization in our synthetic data, which diverges from the input model at m (Section 3). Additionally, because the NIRISS bandpass captures only the red wing of the Na I doublet, the sodium abundance is degenerate with the scattering slope, making it sensitive to both the noise realization and the assumed aerosol properties (Taylor et al., 2023).
@@
Despite its presence in the synthetic data, CO is not strongly detected (1.1 in the traditional case) and our retrievals can only infer an upper limit, likely due to the lack of a strong absorption feature in the NIRISS bandpass (Figure 3). Additionally, the abundance of Na is particularly susceptible to bias due to the lack of a complete absorption feature captured by the NIRISS bandpass; only the red tail of the Na I doublet is captured, which leads to a degeneracy with parameters of the aerosol scattering slope (Taylor et al., 2023).
@@
We assess the performance of these different models with a model comparison based on the Bayes factor and its conversion to a ‘detection significance’ (e.g., model preference, Benneke & Seager, 2013; Welbanks & Madhusudhan, 2021). We find that the global-only and local-only GP frameworks are preferred over the traditional framework by and , respectively. Likewise, the combined local and global GP is preferred over the traditional framework by . All cases strongly indicate the need for a GP, with a moderate preference for a global kernel but no preference regarding a local kernel. We attribute this to the small amplitude of the injected feature, allowing a global GP to account for it without the need for a local kernel (Figure 4). We consider a case with an higher-amplitude injected feature (Appendix A), and find that in the case of the larger feature, the combined global and local GP is preferred over a global GP by ; this implies that for larger features, global kernels will not be enough to marginalize over localized effects, and that the need for a local kernel is partially correlated with the amplitude of the data-model mismatch.
@@
Additionally, we explore whether a GP will artificially inflate noise for data without underlying correlation, providing a false positive. To test this, we generate identical data to that in Section 3, but skip the additional injection of correlated noise, so that the white noise encodes the full underlying uncertainty in the data. We find that the inclusion of any combination of GP kernels does not significantly impact the retrieved parameters in this scenario; additionally, the traditional framework is preferred over all GP frameworks by 3, implying that the GP is not necessary for fitting this dataset accurately. We therefore conclude that the use of a GP-aided retrieval is unlikely to artificially degrade the science output from observations, even in the best-case scenario where a GP is not needed to fully explain the data.

### v2

@@ Abstract
Atmospheric retrievals are essential tools for interpreting exoplanet transmission and eclipse spectra, enabling quantitative constraints on the chemical composition, aerosol properties, and thermal structure of planetary atmospheres. The James Webb Space Telescope (JWST) offers unprecedented spectral precision, resolution, and wavelength coverage, unlocking transformative insights into the formation, evolution, climate, and potential habitability of planetary systems. However, this opportunity is accompanied by challenges: modeling assumptions and unaccounted-for noise or signal sources can bias retrieval outcomes and their interpretation. To address these limitations, we introduce a Gaussian Process (GP)-aided atmospheric retrieval framework that flexibly accounts for unmodeled features and correlated noise in exoplanet spectra. We validate this method on synthetic JWST observations and show that GP-aided retrievals reduce bias in inferred abundances and better capture model-data mismatches than traditional approaches. We also introduce the concept of mean squared error to quantify the trade-off between bias and variance, arguing that this metric more accurately reflects retrieval performance than bias alone. We then reanalyze the NIRISS/SOSS JWST transmission spectrum of WASP-96 b, finding that GP-aided retrievals yield broader constraints on CO2 and H2O, possibly alleviating tension between previous retrieval results and equilibrium predictions. Our GP framework provides precise and accurate constraints while highlighting regions where models fail to explain the data. As JWST matures and future facilities come online, a deeper understanding of the limitations of both data and models will be essential, and GP-enabled retrievals like the one presented here offer a principled path forward.
@@ 2.1 Forward Model
The vertical temperature structure is parameterized following the prescription of Madhusudhan & Seager (2009a). Inhomogeneous clouds and hazes (aerosols) are modeled using an adaptation of the single-sector model from Welbanks & Madhusudhan (2021) where instead of a cloud-top pressure, we consider a grey cloud with an opacity of . The contribution from hazes is modeled as a modification to Rayleigh-scattering with a slope of and enhancement . The inhomogeneous cloud/haze cover is modeled as a linear combination of a clear and cloudy/hazy atmosphere following Line & Parmentier (2016) where is a free parameter describing the cloud/haze fraction. Finally, we include a scaling parameter, R, that accounts for uncertainty in the reported radius of the planet and its degeneracy with the reference pressure (Welbanks & Madhusudhan, 2019). The spectral model is evaluated at a resolution of before being binned to the resolution of the observations. The Bayesian inference is performed using the MultiNest nested sampling algorithm (Feroz et al., 2009) via PyMultiNest (Buchner et al., 2014).
@@ 2.2 Retrieval Framework
where n is the number of wavelength bins, is the observed flux (or transit depth) in the bin with associated uncertainty , and is the model prediction at that wavelength. In this form, the second term of the equation is a constant penalty term representing the white noise of the data.
@@
To account for these correlations, we include a Gaussian process (GP) in our retrieval framework (Figure 1). A GP is defined by two components: a mean function, , which in our case corresponds to the modeled atmospheric spectrum, and a covariance matrix, , which captures correlations between any two input points and is wholly defined by one or more parametric kernel functions that are defined over an infinite domain. These kernels are each defined by their associated hyperparameters (Rasmussen & Williams, 2005).
Each element in the matrix quantifies the covariance between data points and , while the diagonal elements additionally represent the white noise variance at each wavelength bin. When noise is uncorrelated and well-characterized, the off-diagonal terms vanish, and the standard Gaussian likelihood is recovered. A global kernel captures long-range correlations across the entire spectrum, whereas local kernels can be used to model confined regions exhibiting strongly correlated residuals. Figure 2 shows a visual example of a covariance matrix that includes both global and local kernels.
@@
where is the residual between the observed spectrum and the binned model evaluation. is the GP covariance matrix, which consists of the covariance kernel evaluation at each pair of observed wavelengths and . In the traditional retrieval case, where data points are assumed to be uncorrelated, reduces to a diagonal matrix with elements (Figure 2). Under this assumption, Equation 3 simplifies to the standard one-dimensional form shown in Equation 2. Because the GP acts directly on the residuals, it is agnostic to whether discrepancies arise from imperfections in the data or from inadequacies in the model. This allows the retrieval to account for correlated structure in the residuals without requiring explicit identification of its source.
Here, we adopt a covariance matrix composed of both global and local kernels. For the global kernel, we assume a single amplitude and correlation length scale applied uniformly across the spectrum to model correlated noise in the data. Local kernels, in contrast, are fitted to capture correlated structure confined to specific regions. These localized deviations can arise from unmodeled spectral features such as missing absorbers or contamination in the data (e.g., Czekala et al., 2015), as well as from localized artifacts introduced during the data reduction process (e.g., Radica et al., 2023; Holmberg & Madhusudhan, 2023). Accounting for these features with local kernels allows the retrieval to marginalize over regional model–data mismatches that might otherwise bias inferred atmospheric properties. Notably, such biases may not be evident in traditional retrieval frameworks, as residuals in one part of the spectrum can still bias the retrieval outcome for absorbers with features elsewhere in the spectrum (see e.g., Welbanks et al., 2023). In addition, the retrieved hyperparameters of the local kernels can help identify which regions of the spectrum contribute most to poor fits and quantify the extent of the correlated deviations.
@@
where represents the traditional diagonal covariance matrix, is the global kernel, and is the local kernel. The combined kernel, , is evaluated at each pair of observed wavelengths to form the covariance matrix of the data.
@@ 2.3 Kernel Selection
The global kernel is defined by a characteristic amplitude and length scale , where the length scale represents the distance over which wavelength bins exhibit non-negligible correlation (Rasmussen & Williams, 2005). Typically, a positive semi-definite kernel, i.e., one for which for all , is applied across the entire mean function. In this work, we model the global covariance using a squared exponential (SE) kernel,
where is the distance between any two wavelength bins and . This kernel has been used across astrophysical literature, including in Gibson et al. (2012) and Rajpaul et al. (2015). Later studies, including Czekala et al. (2015) and Meech et al. (2022), have used a Matérn-3/2 kernel, which can be considered a rougher variant of the SE kernel. We compare the two on the synthetic spectrum in Section 3 and find little difference in performance, as both are able to reliably constrain the injected correlated noise and provide accurate inferences on absorber abundances. In this work, we focus specifically on observations obtained with NIRISS/SOSS, which has a point spread function (PSF) oversampled by 2–3 pixels along the dispersion axis (Albert et al., 2023). Under these conditions, a Gaussian approximation is well-suited to model the covariance between adjacent pixels. We therefore choose to adopt the SE kernel throughout this work for consistency between kernels.
In addition to a global covariance, localized regions of high residuals can be modeled using local kernels. We assume the local kernel to be a Gaussian function (SE kernel) with amplitude , length scale , and central wavelength , such that
The SE kernel acts as a smooth, low-resolution approximation to unresolved or unmodeled spectral features. While it can effectively account for unknown absorbers or contamination, its functional form is not flexible enough to mimic high-resolution line shapes. This makes it unlikely to compete with physically motivated opacity sources when included in the model, ensuring that the GP captures residual structure without affecting the detection or characterization of real atmospheric features.
@@ 3 Validation on Synthetic NIRISS/SOSS Spectra
The resulting atmospheric model was then binned to the wavelengths and widths from Radica et al. (2023). In addition to white noise drawn from the reported precision of the JWST/NIRISS SOSS spectrum in Radica et al. (2023), we inject global correlated noise across the entire wavelength range, with an amplitude of ppm and a correlation length scale of mẆe also inject a localized high-correlation feature centered at 1.65 m. The final synthetic spectrum is generated as a single random draw from a multivariate normal distribution, incorporating both white and correlated noise. This approach simulates realistic observational scatter while preserving the imposed correlation structure. In doing so, we test whether the retrieved hyperparameters can inform inferences about the underlying correlation structure and the specific wavelengths most affected. The model, injected noise, and resulting synthetic data are shown in Figure 3. In addition to the noticeable feature at 1.65 m, the simulated noise results in a generally higher transit depth relative to the input model at wavelengths m.
@@
which corresponds to the specific case of Equation 4 with one local kernel.
Table 1 provides the priors for the parameters and hyperparameters considered. We elect to sample along uniform priors for each hyperparameter, rather than log-uniform priors, as this provides the most informative hyperparameter posterior distributions. We motivate this choice further in Section 5.1. However, we caution that prior selection should be done on a case-by-case basis, and that improper choice of prior can lead to biasing effects in the posterior distribution, particularly when the expected values for a parameter cover multiple orders of magnitude.
@@ 3.1 Retrieval Performance on Synthetic WASP-96 b Data
Although the traditional retrieval produces the most precise posterior distributions, it yields highly inaccurate results due to the injected correlated noise. The median retrieved abundances of H2O, CO2, and CH4 are offset from the true values by 1.24, 1.89, and 4.64 dex, respectively (Figure 4), demonstrating general overestimates due to unmodeled noise effects, particularly for CH4 abundance.
@@
In the “local-only” retrieval, a single local kernel is used to capture a high-correlation feature in the spectrum. This configuration enables the retrieval to constrain the injected feature’s amplitude and central wavelength, but it does not account for the global correlations present across the spectrum. As a result, the retrieval remains biased in key atmospheric parameters: the H2O and CO2 abundances are still offset by 0.62 and 0.81 dex, respectively. However, the local kernel successfully mitigates the effect of the injected feature near 1.7 m, which overlaps with a CH4 absorption band, resulting in a CH4 abundance much closer to the true value than in the traditional case (0.89 dex offset rather than 4.64 dex). This suggests that while local kernels can improve model–data agreement in targeted regions, they are insufficient for capturing broader correlated structure.
@@
In contrast to the traditional case, the retrieval using a global kernel recovers all gas abundances to within 2 of their true values, but at the expense of precision. For example, the CO2 abundance is now much more accurate (0.36 dex offset compared to 1.89 dex in the traditional case), but the 1 posterior interval spans nearly 4 dex (Figure 4). As a result, the retrieval provides only weak constraints that are insufficient to inform parameters such as atmospheric metallicity or the carbon-to-oxygen ratio.
@@
Finally, we consider the retrieval that includes both global and local kernels. This configuration enables the GP to accurately constrain all hyperparameters (Figure 4, bottom), including successful identification of the injected feature at 1.65 m. It also yields atmospheric parameters that are much closer to the true values, without substantially increasing the posterior variance. The retrieved median abundances of H2O, CO2, and CH4 differ from their true values by only 0.19, 0.37, and 0.20 dex, respectively (Figure 4). The Na abundance is measured with higher precision than in the global-only case and greater accuracy than in the traditional and local-only retrievals, although it is still overestimated relative to the input. This overestimation may be partially attributable to the specific noise realization in our synthetic data, which diverges from the input model at m (Section 3). Additionally, because the NIRISS bandpass captures only the red wing of the Na I doublet, the sodium abundance is degenerate with the scattering slope, making it sensitive to both the noise realization and the assumed aerosol properties (Taylor et al., 2023).
@@
Despite its presence in the synthetic data, CO is not significantly detected (a detection significance, e.g., Trotta (2008), of 1.1 in the traditional case) and our retrievals can only infer an upper limit, likely due to the lack of a strong absorption feature in the NIRISS bandpass (Figure 3). Additionally, the abundance of Na is particularly susceptible to bias due to the lack of a complete absorption feature captured by the NIRISS bandpass; only the red tail of the Na I doublet is captured, which leads to a degeneracy with parameters of the aerosol scattering slope (Taylor et al., 2023).
@@
We assess the performance of these different models with a model comparison based on the Bayes factor and its conversion to a ‘detection significance’ (e.g., model preference, Trotta, 2008; Benneke & Seager, 2013; Welbanks & Madhusudhan, 2021). We find that the global-only and local-only GP frameworks are preferred over the traditional framework by (log-Bayes factor ) and () respectively. Likewise, the combined local and global GP is preferred over the traditional framework by (). All cases correspond to a strong indication of the need for a GP (Trotta, 2008) but no statistically significant preference regarding the use of a local kernel when a global kernel is also used (, ). We attribute this to the small amplitude of the injected feature, allowing a global GP to account for it without the need for a local kernel (Figure 4). We consider a case with an higher-amplitude injected feature (Appendix A), and find that in the case of the larger feature, the combined global and local GP is preferred over a global GP by (); this implies that for larger features, global kernels will not be enough to marginalize over localized effects, and that the need for a local kernel is partially correlated with the amplitude of the data-model mismatch.
@@
Additionally, we explore whether a GP will artificially inflate noise for data without underlying correlation, providing a false positive. To test this, we generate identical data to that in Section 3, but do not add correlated noise, so that the white noise encodes the full underlying uncertainty in the data. We find that the inclusion of any combination of GP kernels does not significantly impact the retrieved parameters in this scenario; additionally, the traditional framework is preferred over all GP frameworks by 3, implying that the GP is not necessary for fitting this dataset accurately. We therefore conclude that our use of a GP-aided retrieval is unlikely to artificially degrade the science output from observations in this work, even in the best-case scenario where a GP is not needed to fully explain the data. This is, however, dependent on both the GP kernel parameterization and expected posterior distribution precisions given a white noise model, among other effects. It should therefore be considered for each case independently. A more complex GP model is more likely to degrade the science output for a given dataset, as model flexibility increases at the cost of precision.
The Bayes factor is used here to indicate the strength of the preference for the retrieval that includes a GP noise model rather than the traditional white noise model. However, it is important to note that by changing the noise model to be more flexible, the GP noise model case can actually be considered a less complex noise model, allowing for structured correlations in the data to be considered within the noise model rather than being attributed to a more complex physical model. The Bayes factor is also highly dependent on the selection of parameter and hyperparameter prior ranges, as well as the assumption that the models being compared are both approximately valid interpretations of the data. Here, we select hyperparameter priors that we expect to be broad enough to not bias the hyperparameters towards a given value, while also remaining within physical bounds (i.e., avoiding correlation length scales larger than the wavelength range of the data or unphysical kernel amplitudes for a transmission spectrum). Bayes factors should therefore not be considered definitive indicators of a need for a GP, and are rather a conditional quantification of the comparison between the GP and non-GP noise models combined with the atmospheric models used in this work. We include further discussion of the selection of hyperparameter priors and their effects in Section 5.1.